### 20 Mar 2026 v1

# From Plausibility to Verifiability: Risk-Controlled Generative OCR for Vision-Language Models

Abstract.
Modern vision-language models (VLMs) can act as generative OCR engines, yet open-ended decoding can expose rare but consequential failures. We identify a core deployment misalignment in generative OCR. Autoregressive decoding favors semantic plausibility, whereas OCR requires outputs that are visually grounded and geometrically verifiable. This mismatch produces severe errors, especially over-generation and unsupported substitutions, creating deployment risk even when benchmark accuracy remains high. We therefore formulate frozen VLM OCR as a selective accept/abstain problem and propose a model-agnostic Geometric Risk Controller. The controller probes multiple structured views of the same input, applies lightweight structural screening, and accepts a transcription only when cross-view consensus and stability satisfy predefined criteria, yielding a small family of operating points. Experiments on frozen VLM backbones and standard OCR benchmarks show consistent reductions in extreme-error risk and catastrophic over-generation at predictable coverage costs. Reliable deployment of generative OCR with frozen VLMs benefits from explicit system-level risk control rather than unconstrained generation.
copyright: none
1. Introduction
Figure 1. System overview of the proposed Geometric Risk Controller (GRC). Given multi-view queries to a frozen vision-language model, the controller applies structural screening and cross-view consensus to produce a candidate string $s^{*}$, where $s^{*}$ is the consensus transcription across views. The operating point then decides whether to accept $s^{*}$ or abstain.
System overview of the Geometric Risk Controller for deployment-controlled generative OCR. An input crop is expanded into K geometric views, which are queried with a frozen vision-language model backbone. The returned transcriptions enter a controller layer with preprocessing, structural screening, cross-view consensus, and an operating-point controller. The final decision is accept or abstain, yielding either an accepted transcription or no output.
Optical character recognition (OCR) is increasingly treated as an emergent capability of modern vision-language models (VLMs) (Lv et al., 2024; Yu et al., 2024). In practice, this has encouraged the use of frozen VLMs as direct generative OCR engines. However, such usage remains fragile in deployment. Under open-ended autoregressive decoding, a model may generate continuations that extend beyond the visible text or produce semantically plausible yet visually unsupported substitutions. These failure modes violate a key requirement for user-facing OCR, namely geometric verifiability, meaning that the emitted text should be locally supportable by image evidence under geometric constraints.
Traditionally, OCR systems are evaluated by average-case accuracy, namely whether the predicted string matches the ground-truth transcription on a benchmark (Karatzas et al., 2013). Here the focus shifts to deployment risk, defined by the risk exposed on the outputs returned to users under a fixed inference protocol. Under this view, generative OCR with frozen VLMs becomes a selective prediction problem with abstention, where the system either emits a transcription or abstains, and abstention is part of the deployment contract rather than a convenience feature (Geifman and El-Yaniv, 2017; Wen et al., 2025). Making the accept-versus-abstain decision explicit turns evaluation from a single average-case score into an auditable risk–coverage trade-off under fixed protocols (Geifman and El-Yaniv, 2017; Angelopoulos et al., 2024). The most harmful OCR failures are both rare and highly sensitive to the inference protocol, especially the prompt template and decoding strategy. Open-ended generative decoding is optimized for coherent continuation, so when visual evidence is weak or ambiguous, the model may fall back on language priors and produce semantically plausible but visually unsupported outputs (Li et al., 2023). Consequently, stronger average benchmark performance does not necessarily reduce the severe-error tail that drives user-facing exposure. Reliability also should not depend on ad hoc prompting, since even minor prompt changes can materially alter model behavior (Zhao et al., 2021; Rudman et al., 2026). Decoding-time heuristics can likewise change hallucination frequency, making behavior harder to reproduce and audit even under a nominally fixed deployment setup (Huang et al., 2024b). Our target setting is deployment-driven. Foundation models are often shared across tasks and products, making per-application retraining for OCR costly or impractical. We therefore seek an external control layer that operates at inference time and reduces exposed risk without modifying model internals. This layer is not an after-the-fact heuristic filter; rather, the accept/abstain rule is specified as part of the deployment contract. We therefore propose a deployment-centric risk-control framework for generative OCR with frozen VLMs. At its core is the Geometric Risk Controller (GRC), which converts open-ended decoding into a selective system under a fixed accept/abstain contract. Under a fixed multi-view protocol with $K$ geometrically related views, GRC probes the frozen model through controlled input variations, applies lightweight structural validity checks, and accepts a transcription only when cross-view agreement is sufficiently stable. A single strictness knob $m$ exposes a small family of operating points on a risk–coverage curve, enabling explicit operating-point control rather than ad hoc retuning (Geifman and El-Yaniv, 2017; Angelopoulos et al., 2024). We provide full protocol details and component ablations, and report results at fixed operating points so that trade-offs remain comparable across models and datasets. While related in spirit to recent test-time scaling strategies for compact VLMs based on augmentation and aggregation (Kaya et al., 2026), our formulation is specialized to OCR failure modes and instantiated as a fixed, deployment-auditable accept/abstain contract. Our contributions are threefold:
(1) Reframe. We recast frozen VLM-based OCR as a deployment-control problem, introducing geometric verifiability and deployment-oriented risk primitives that expose long-tail and catastrophic failures beyond mean accuracy.
(2) Control. We propose a model-agnostic Geometric Risk Controller that converts open-ended generative OCR into a fixed-protocol selective system with an explicit strictness knob.
(3) Validate. Across multiple frozen VLM backbones and scene-text benchmarks, we show that the controller consistently suppresses exposed long-tail risk while offering controllable coverage trade-offs; a compact query-budget analysis further illustrates the associated inference-cost trade-off.
2. Related Work
We organize prior work around deployment reliability in frozen VLM OCR, focusing on why stronger OCR capability alone does not guarantee reliable deployment, how prior work introduces selective accept/abstain interfaces to address this gap, and what forms of test-time evidence are needed to make such interfaces auditable under a fixed inference protocol.
2.1. Generative OCR Backbones and the Deployment Misalignment
Frozen VLMs as Generative OCR Engines. Vision-language models (VLMs) are increasingly used as OCR engines (Kim et al., 2022). Systems such as KOSMOS-2.5 (Lv et al., 2024), TextHawk (Yu et al., 2024), and OCEAN-OCR (Chen et al., 2025) show strong end-to-end transcription ability across several settings. However, stronger OCR capability does not by itself guarantee deployment reliability. Under open-ended generation, autoregressive decoding optimizes next-token likelihood rather than explicit visual verifiability, leading to unsupported substitutions and runaway continuations (Huang et al., 2024b). These failures are sensitive to prompting and decoding choices (Rudman et al., 2026; Schmalfuss et al., 2025; Huang et al., 2024b), complicating auditing and control in deployment. Recent analyses further link such behaviors to broader visual-grounding weaknesses in multimodal LLMs (Tong et al., 2024).
Average-Case Metrics versus Deployment Exposure. Benchmarks such as the ICDAR series (Karatzas et al., 2013, 2015; Huang et al., 2019), OCRBench (Liu et al., 2024; Fu et al., 2025), and SVT (Wang et al., 2011) emphasize aggregated scores such as recognition accuracy, CER, and exact match. While these metrics capture average capability, they can understate rare but high-cost tail failures (Amit et al., 2025). For generative OCR, this creates a deployment misalignment between benchmark capability and exposed-output reliability, motivating an explicit accept–abstain formulation (Mao et al., 2025) under selective prediction and risk–coverage principles (Geifman and El-Yaniv, 2017).
2.2. Selective Deployment and Risk-Controlled Decision Interfaces
Selective deployment requires an explicit accept–abstain interface, not merely a stronger recognizer. Selective prediction formalizes this interface through the coverage–risk trade-off (Geifman and El-Yaniv, 2017; Wen et al., 2025; Chow, 1970), while Conformal Risk Control (CRC) formalizes finite-sample control of expected risk under a specified loss (Wang et al., 2025; Angelopoulos et al., 2024; Angelopoulos and Bates, 2021). In generative systems, however, abstention is often based on internal confidence or self-evaluation (Ren et al., 2023; Mao et al., 2025), and such signals can be prompt- and protocol-sensitive (Schmalfuss et al., 2025) while remaining overconfident on hallucinated outputs (Zollicoffer et al., 2025; Mao et al., 2025; Woo et al., 2025). Our controller adopts the same deployment-level framing but remains empirical rather than guarantee-providing, basing abstention on protocol-induced exogenous evidence—cross-view consistency and structural stability—to define auditable fixed-protocol operating points.
2.3. Verification Signals under Fixed Inference Protocols
Test-time scaling methods such as test-time augmentation, multi-crop inference, and self-consistency are widely used to improve performance and robustness through aggregation, reranking, or voting (Snell et al., 2024; Lee et al., 2025; Wang et al., 2023; Kaya et al., 2026; Chou et al., 2025a, b). Our setting uses repeated inference for a different purpose. Under a fixed protocol, cross-view agreement and stability serve as decision evidence for accept–abstain control rather than as heuristics for improving accuracy (Khan and Fu, 2024). Related OCR and structured extraction systems also use lightweight output constraints, such as character-set, length, and format checks, to suppress obviously invalid generations (Willard and Louf, 2023), but these string-level filters do not directly establish visual support (Li et al., 2025) and can remain misleading when the model is consistently wrong (Huang et al., 2024a). We therefore treat cross-view consistency under controlled geometric perturbations as exogenous evidence of visual support, turning test-time consistency into a fixed-protocol control signal for auditable accept–abstain decisions.
3. Methods
3.1. Setting and Control Contract
We study generative OCR with a frozen vision-language model (VLM) used as a black-box generator. Given an image crop $I\in\mathcal{I}$, a fixed prompt template $p$, and a fixed deterministic decoding configuration $\delta$, the frozen backbone returns
$y=G_{\delta}(I,p;\theta)\in\mathcal{Y},\qquad\theta\ \text{frozen}.$
We treat $G_{\delta}$ as the black-box mapping induced by the frozen backbone under the fixed prompt-and-decoding setup; model internals are not used, and control relies only on queried image views and returned strings.
Our goal is not to modify the backbone, but to place it inside a selective accept–abstain contract. Accordingly, the deployed OCR system is
$f:\mathcal{I}\to\mathcal{Y}\cup\{\bot\},$
where $\bot$ denotes abstention. The map $f$ is induced by an inference-time evidence protocol $\pi$ and an acceptance rule applied to that evidence.
Specifically, a protocol $\pi$ extracts from the input crop a candidate transcription $\hat{y}(I)\in\mathcal{Y}$ together with an evidence summary $e(I)$. An acceptance rule $A$ then determines whether the candidate is exposed:
$f_{\pi}(I)=\begin{cases}\hat{y}(I),&\text{if }A\!\big(e(I)\big)=1,\\ \bot,&\text{otherwise.}\end{cases}$
Conventional frozen-VLM OCR is recovered as the degenerate always-accept case, in which the protocol reduces to a single standard query and the returned output is always exposed.
3.2. Evidence Protocol: Multi-view Geometric Probing
We instantiate $\pi$ through multi-view probing of the same input crop. Let $\mathcal{T}$ denote a family of admissible image transforms that apply mild geometric perturbations, such as small translations, crop jitter, or scale variation, while preserving nominal text content.
A fixed protocol $\pi$ specifies the transform family $\mathcal{T}$, a view-selection rule, and the total number of queried views $K$. The original crop is always included as the anchor view,
$T_{1}=\mathrm{Id},\qquad I_{1}=I.$
The remaining views are
$I_{k}=T_{k}(I),\qquad k=2,\dots,K,$
where the selection rule for $\{T_{k}\}_{k=2}^{K}$ is part of the protocol. In our implementation, the non-anchor transforms are sampled from a fixed distribution over admissible transforms. Each view is queried with the same prompt template and decoding configuration as in Sec. 3.1:
$y_{k}=G_{\delta}(I_{k},p;\theta),\qquad k=1,\dots,K.$
This yields the evidence record
$R_{\pi}(I)=\{(I_{k},y_{k})\}_{k=1}^{K}.$
Under the fixed protocol, we treat cross-view stability as an operational proxy for geometric support and cross-view disagreement as a risk signal. These exogenous signals are then used for accept–abstain control.
Figure 2. Representative accept–abstain evidence patterns under the fixed K=5K{=}5 protocol at the default operating point (m=3m{=}3). Fragmented valid outputs cause abstention, a majority can still be rejected under high dispersion, and borderline consensus is accepted when the valid views remain tightly clustered.
A single-column figure with three rows showing evidence patterns under a fixed five-view protocol. In the first row, all five views are valid but disagree, yielding low exact consensus and abstention. In the second row, four views are valid and one is invalid; although a majority mode exists, the remaining valid outputs are highly dispersed, so the sample is rejected. In the third row, all five views are valid and most cluster around a dominant mode, so a borderline vote is still accepted. Each row includes a summary box with the number of valid views, the selected mode, the vote fraction q, the dispersion Delta, and the final decision.
3.3. Canonicalization and Structural Screening
To make cross-view evidence comparable under the fixed protocol, we first map each returned string to a canonical form
$\tilde{y}_{k}=\mathrm{Norm}(y_{k})\in\mathcal{Y},$
where $\mathrm{Norm}(\cdot)$ is a protocol-fixed normalization rule that removes only presentation variance, such as whitespace normalization and, when specified by the task protocol, case canonicalization, without changing the underlying transcription content.
We then apply a structural screening rule
$v_{k}=\mathbb{I}[\tilde{y}_{k}\in\mathcal{V}]\in\{0,1\},$
where $\mathcal{V}\subseteq\mathcal{Y}$ is a protocol-fixed set of lightweight admissibility constraints. These constraints are label-agnostic and serve only to reject degenerate continuations, not to recognize text or estimate confidence.
The main admissibility constraint in $\mathcal{V}$ is a conservative geometric length bound. For each view $I_{k}$, we compute an upper bound $L_{\mathrm{geom}}(I_{k})$ from foreground geometry after scale normalization and dominant-axis span estimation. A decoded string is marked invalid if
$|\tilde{y}_{k}|>L_{\mathrm{geom}}(I_{k}).$
This bound is not intended to estimate the exact ground-truth length; it only suppresses implausibly overlong continuations.
The resulting processed evidence is
$\widetilde{R}_{\pi}(I)=\{(\tilde{y}_{k},v_{k})\}_{k=1}^{K}.$
Let
$\Omega(I)=\{k:v_{k}=1\},\qquad n(I)=|\Omega(I)|.$
All downstream evidence statistics are computed only over $\Omega(I)$. This induces a minimum-evidence gate with threshold $K_{\min}$.
3.4. Decision Functional: Consensus, Stability, and Operating-Point Control
For samples with $n(I)\geq K_{\min}$, let $s^{*}(I)$ denote the unique mode of the valid canonicalized outputs $\{\tilde{y}_{k}:k\in\Omega(I)\}$ when it exists. If the mode is not unique, the system abstains. Otherwise, the candidate transcription is
$\hat{y}_{\pi}(I)=s^{*}(I),$
with vote fraction
$q(I)=\frac{1}{n(I)}\sum_{k\in\Omega(I)}\mathbb{I}\!\left[\tilde{y}_{k}=s^{*}(I)\right]\in[0,1].$
To measure residual disagreement around the candidate, we use the bounded normalized edit distance
$d(a,b)=\min\!\left(1,\frac{\mathrm{ED}(a,b)}{\max(1,\max(|a|,|b|))}\right)\in[0,1],$
where $\mathrm{ED}(\cdot,\cdot)$ is Levenshtein edit distance, and define the candidate-centered dispersion
$\Delta(I)=\frac{1}{n(I)}\sum_{k\in\Omega(I)}d\!\left(\tilde{y}_{k},s^{*}(I)\right).$
The resulting evidence summary is
$e_{\pi}(I)=\big(n(I),\,q(I),\,\Delta(I)\big).$
The controller uses a consensus threshold $\tau(m)$ and a fixed dispersion threshold $\kappa$. In the main operating-point family, the strictness index $m$ affects only the minimum required vote fraction, with $\tau(m)$ nondecreasing in $m$, while $\kappa$ is held fixed. Larger $m$ therefore induces a stricter acceptance region through a stronger consensus requirement under the common stability gate $\Delta(I)\leq\kappa$.
The instantiated controller is
[…]
Algorithm 1 Multi-view geometric probing and accept–abstain control
1:Crop II, frozen generator $G_{\delta}$, prompt $p$, protocol/controller instantiation $(\pi,m)$
2:Output transcription $s^{*}$ or abstention $\bot$
[…]
4:for $k=1,\dots,K$ do
5:  $\tilde{y}_{k}\leftarrow\mathrm{Norm}(G_{\delta}(I_{k},p;\theta))$,  $v_{k}\leftarrow\mathbb{I}[\tilde{y}_{k}\in\mathcal{V}]$
[…]
7:if $|\Omega|<K_{\min}$ then
8:  return $\bot$
9:Determine the unique mode $s^{*}$ of $\{\tilde{y}_{k}:k\in\Omega\}$
10:if no unique mode exists then
11:  return $\bot$
[…]
13:if $q\geq\tau(m)$ and $\Delta\leq\kappa$ then
14:  return $s^{*}$
15:return $\bot$
3.5. Coverage, Risk, and Operating Points
Selective OCR is evaluated jointly by coverage and conditional risk on exposed outputs. For the controlled system $f_{\pi,m}(I)\in\mathcal{Y}\cup\{\bot\}$, coverage is
$\mathrm{Cov}(\pi,m)=\Pr\!\left[f_{\pi,m}(I)\neq\bot\right].$
If the protocol uses randomized view selection, the probability is taken jointly over the data distribution and the induced protocol randomness.
All controlled-system risk metrics are computed on the covered subset. We report mean CER and upper-tail CER (P99) on covered outputs, and define the conditional catastrophic exposure rate as
$\mathrm{Meltdown}@\delta(\pi,m)=\Pr\!\left[\mathrm{CER}\!\left(f_{\pi,m}(I),y^{\mathrm{gt}}\right)\geq\delta\ \middle|\ f_{\pi,m}(I)\neq\bot\right],$
with $\delta=2$ in our main setting. Since CER is normalized by ground-truth length, severe over-generation may yield $\mathrm{CER}>1$.
We report results on a small set of predefined operating points $m$ under a common protocol $\pi$. In the main instantiation used in this paper, these operating points correspond to increasing consensus thresholds $\tau(m)$ while the dispersion threshold $\kappa$ is fixed. These settings are fixed before evaluation and shared across models and benchmarks.
For comparison, we report baseline risk on the full test set and controlled-system risk on the covered subset together with coverage. As an additional reference point, we include an approximately coverage-matched external selective baseline in Sec. 4.3.
4. Experiments
4.1. Experimental Setup
Backbones and datasets. We evaluate the controller on three frozen vision-language backbones used as black-box generators: LLaVA-Phi3 (3.8B) (Zhu et al., 2024), Gemma3 (4B) (Team et al., 2025), and GLM-OCR. Experiments use the official test sets of IIIT5K (Mishra et al., 2012) and ICDAR 2013 (Karatzas et al., 2013).
Protocol instantiation. Unless otherwise noted, the common protocol uses $K=5$ geometrically related views, consisting of one anchor view and four additional views sampled from a fixed admissible transform distribution over translation, crop jitter, and scale variation. Returned strings are canonicalized by $\mathrm{Norm}(\cdot)$ and filtered by the admissibility screen $\mathcal{V}$, with $K_{\min}=3$.
Operating points. We use predefined strictness levels $m\in\{1,3,5\}$. In the main instantiation, these correspond to increasing consensus thresholds
$\tau(1)=0.1,\qquad\tau(3)=0.5,\qquad\tau(5)=0.9,$
while the dispersion threshold is fixed at $\kappa=0.4$. All settings are fixed before evaluation and shared across models and benchmarks. Unless otherwise stated, the main results use $m=3$.
Metrics and reporting protocol. Following Sec. 3.5, we report coverage together with covered-subset risk, including mean CER, upper-tail CER (P99), and $\mathrm{Meltdown}@2$. Since the baseline never abstains, its risk is reported on the full test set, whereas controlled-system risk is reported on covered outputs. Whenever the controlled system applies output canonicalization, the same canonicalization is also applied to the baseline. Sec. 4.3 further includes an approximately coverage-matched external baseline.
4.2. Main Results
Figure 3. Qualitative OCR cases under the fixed deployment protocol. Each panel shows the crop, ground truth, the always-accept baseline, and the GRC decision at the default operating point ($m=3$). Cases include over-generation, unsupported substitution, correct acceptance, cross-view instability, a stable-but-wrong residual failure, and a borderline accepted case. The evidence tag reports valid-view count, consensus fraction $q$, and geometric length estimate $L_{\mathrm{geom}}$; for the instability case, it reports dispersion $\Delta$ instead of $L_{\mathrm{geom}}$.
[…]
Table 1. Main deployment-level results at the fixed operating point ($K=5$, $m=3$).
Dataset Metric Model
LLaVA-Phi3 Gemma3 GLM-OCR
[…]
Coverage is reported for the controlled system only. For risk metrics, each entry is shown as Baseline/System, where Baseline is computed on the full test set and System on the covered subset. CER is reported in %; MD@2 in ‰.
Table 1 reports the main deployment-level results under the common fixed protocol with $K=5$ at the moderate operating point $m=3$. These results address the deployment setting introduced in Sec. 1 by quantifying the error risk exposed on the outputs that the system actually returns under a fixed selective contract.
Across all $3\times 2$ model–dataset settings, the Geometric Risk Controller maintains high coverage while substantially reducing exposed risk relative to the always-accept baseline. Relative to baseline full-set exposure, the controlled system consistently lowers mean CER, suppresses upper-tail error, and reduces catastrophic exposure as measured by $\mathrm{Meltdown}@2$. These results align with the intended deployment role of the controller, turning open-ended generative OCR into a selective system that trades a moderate amount of coverage for a substantially safer set of exposed outputs. Because the baseline never abstains, its risk is reported on the full test set; for the controlled system, risk is reported on the covered subset together with coverage. Figure 3 shows representative qualitative outcomes under the fixed protocol, including over-generation, unsupported substitution, instability-driven abstention, and stable-but-wrong residual failures.
4.3. Approximately Coverage-Matched External Baseline Comparison
We next test whether the gains of GRC can be explained by standard confidence-based abstention alone. To this end, we compare against an external confidence-threshold selective baseline (Conf.-Thr.) adapted from the vanilla selective-prediction component of Srinivasan et al. (2024). We retain only the threshold-based abstention rule, so that the comparison isolates a standard internal-confidence selective baseline under the OCR setting.
For fairness, this baseline uses the same frozen backbone, prompt template, deterministic decoding, and output canonicalization as the always-accept OCR baseline. It operates on a single view and accepts a transcription only when the model’s internal confidence exceeds a calibrated threshold. Confidence is instantiated by the mean token log-probability of the generated transcription, and the threshold is calibrated on a held-out split to target the coverage of our default operating point ($m=3$) for each backbone–dataset setting. Realized test-set coverage is therefore approximately, rather than exactly, matched.
Table 2 reports the comparison. At similar coverage, the confidence-threshold baseline reduces some ordinary errors, but remains much weaker at suppressing catastrophic exposure. GRC achieves lower CER mean and lower $\mathrm{Meltdown}@2$ in all six settings, with the largest gains on the more failure-prone LLaVA-Phi3. This suggests that single-view internal confidence does not capture the instability revealed by multi-view external evidence.
Table 2. Approximately coverage-matched comparison with an external confidence-threshold selective baseline adapted from Srinivasan et al. (2024). For each backbone–dataset setting, the threshold is calibrated on a held-out split to target the coverage of our default operating point ($m=3$). Coverage and mean CER are reported in %; Meltdown@2 (MD@2) is reported in ‰.
[…]
IIIT5K LLaVA-Phi3 Conf.-Thr. 90.0 30.9 19.6
Ours 89.5 8.4 0.3
Gemma3 Conf.-Thr. 95.8 7.6 2.2
Ours 94.9 3.7 0.0
GLM-OCR Conf.-Thr. 94.8 4.4 2.2
Ours 95.7 4.0 0.7
ICDAR13 LLaVA-Phi3 Conf.-Thr. 88.2 82.7 31.4
Ours 89.0 9.3 3.7
Gemma3 Conf.-Thr. 95.5 22.5 15.1
Ours 93.0 3.7 1.8
GLM-OCR Conf.-Thr. 95.9 3.2 5.8
Ours 96.2 2.3 0.9
4.4. Operating-Point Control
Figure 4. Risk–coverage trajectories under the common fixed protocol for all three backbones on IIIT5K and ICDAR13. Each point is a predefined operating point indexed by $m$; larger $m$ yields lower coverage and lower covered-output risk. Filled markers denote the default operating point $m=3$.
Risk–coverage plots for three backbones: GLM-OCR, Gemma3, and LLaVA-Phi3. The x-axis is coverage and the y-axis is system mean CER on covered outputs. Each subplot shows two curves, one for IIIT5K and one for ICDAR13, with operating points indexed by m. As m increases, the operating point moves toward lower coverage and lower risk. Filled markers indicate the default operating point m=3.
Figure 5. Component effects and comparison to an external confidence-threshold baseline on LLaVA-Phi3 under the fixed protocol (K=5K{=}5, m=3m{=}3). Bars show Meltdown@2; labels show coverage. Full GRC achieves the lowest catastrophic exposure on both datasets while maintaining competitive coverage.
A single-column figure with two side-by-side bar charts labeled IIIT5K and ICDAR13. Each chart compares five methods—Baseline, Conf-thr, Consensus, Structural, and Full GRC—using Meltdown at 2 on the vertical axis. Numeric labels above the bars report coverage. In both datasets, Full GRC has the lowest catastrophic exposure, while the confidence-threshold baseline reduces exposure relative to the always-accept baseline but remains substantially worse than Full GRC.
A key property of the controller is that the strictness index $m$ does not modify the backbone or protocol; it selects one predefined operating point under a fixed selective contract. In our main instantiation, increasing $m$ raises the required consensus threshold $\tau(m)$ while the dispersion gate remains fixed, making acceptance more conservative and trading coverage for lower exposed risk.
Figure 4 visualizes this behavior across all three backbones on both IIIT5K and ICDAR13. In each case, increasing $m$ moves the system along a consistent risk–coverage frontier toward lower coverage and lower covered-output mean CER. This behavior is consistent with interpreting $m$ as a practical deployment knob, since the operator does not modify the backbone or protocol, but instead selects an operating point from a fixed and auditable family.
4.5. Ablation Study
Table 3. Component ablation under the common fixed protocol ($K=5$, $m=3$). w/o structural removes the protocol-fixed admissibility screen, w/o consensus removes the cross-view consensus control, and Full GRC applies both. Coverage and CER are reported in %. Meltdown@2 is omitted here and shown in Fig. 5.
Method LLaVA-Phi3 Gemma3 GLM-OCR
[…]
IIIT5K
Baseline 100.0 110.5 100.0 7.4 100.0 6.3
w/o structural 95.3 21.8 98.8 4.6 99.6 5.5
w/o consensus 91.8 11.2 95.8 4.3 96.0 4.5
Full GRC 89.5 8.4 94.9 3.7 95.7 4.0
ICDAR13
Baseline 100.0 147.4 100.0 20.1 100.0 3.3
w/o structural 93.1 31.7 96.1 10.4 99.6 3.3
w/o consensus 93.7 17.4 96.0 8.7 96.5 2.5
Full GRC 88.9 9.3 93.0 3.7 96.2 2.3
We next ablate the two decision channels of the controller under the same fixed protocol used above. Table 3 compares the full system with two partial variants. The w/o structural variant removes the protocol-fixed admissibility screen while retaining cross-view consensus control, whereas w/o consensus removes vote-fraction and dispersion control while retaining structural screening.
The two components play distinct and complementary roles. The w/o consensus variant is particularly effective at filtering obviously degenerate continuations and often delivers strong reductions in catastrophic exposure. The w/o structural variant captures cross-view instability and usually improves over the baseline, but its gains are less uniform, especially on already stronger backbones. Neither partial variant matches the full controller.
Full GRC consistently delivers the strongest overall deployment behavior across models and datasets, achieving the lowest exposed risk in every setting while maintaining high coverage. Taken together, these results show that the controller is not a loose collection of heuristics, but a complementary combination of admissibility screening and cross-view consensus control.
4.6. Operating-Point Endpoints and Practical Trade-offs
Table 4. Representative endpoint operating points under the fixed protocol ($K=5$). Coverage (Cov.) is reported in %, and Meltdown@2 (MD@2) is reported in ‰.
IIIT5K ICDAR13
Model Set. Cov. $\uparrow$ MD@2 $\downarrow$ Cov. $\uparrow$ MD@2 $\downarrow$
LLaVA-Phi3 Baseline 100.0 33.7 100.0 35.6
$m=1$ 91.3 0.3 92.1 3.7
$m=5$ 67.7 0.0 68.3 0.9
Gemma3 Baseline 100.0 3.7 100.0 16.4
$m=1$ 95.8 0.0 95.1 1.8
$m=5$ 82.5 0.0 82.0 1.8
GLM-OCR Baseline 100.0 3.7 100.0 1.8
$m=1$ 95.9 1.0 96.5 0.9
$m=5$ 90.5 0.0 91.0 0.9
Figure 4 shows that predefined operating points trace a risk–coverage frontier under a fixed protocol. Because deployment is typically selected from a small set of discrete operating-point options, Table 4 reports three representative settings, namely the always-accept baseline, a permissive setting ($m=1$), and a conservative setting ($m=5$). Stricter operating points reduce catastrophic exposure at the cost of lower coverage. This separation is most pronounced on LLaVA-Phi3 and remains consistent on Gemma3 and GLM-OCR. The default setting $m=3$ used in Sec. 4.2 lies between these endpoints and provides a practical balance between availability and exposed-output reliability.
Table 5 reports a compact query-budget check on LLaVA-Phi3. The single-pass always-accept baseline is included for reference, while $K\in\{3,5,7\}$ varies the evidence budget within the multi-view controller. Reducing to $K=3$ slightly lowers catastrophic exposure but also lowers coverage, while increasing to $K=7$ adds cost with little gain over $K=5$. This supports $K=5$ as a practical default.
Table 5. Compact query-budget check for LLaVA-Phi3 at $m=3$. Relative time is normalized by the single-pass baseline; coverage is reported in %, and Meltdown@2 in ‰.
$K$ Time IIIT5K ICDAR13
Cov. $\uparrow$ MD@2 $\downarrow$ Cov. $\uparrow$ MD@2 $\downarrow$
Baseline 1.0x 100.0 33.7 100.0 35.6
3 2.7x 85.4 0.0 86.3 1.8
5 4.5x 89.5 0.3 88.9 3.7
7 6.3x 89.5 0.3 89.1 3.7
5. Discussion and Limitations
GRC is designed for deployment control rather than backbone improvement. Under a fixed protocol, it reduces unstable user-facing exposure, but it does not eliminate all accepted errors. The main residual failure mode is stable-but-wrong consensus, where multiple admissible views support the same incorrect word and the system still accepts it. This marks a core boundary of the current design. Cross-view agreement provides evidence of stability, but not proof of correctness. As a result, GRC is most effective at filtering unstable and catastrophic exposure, and less effective at detecting errors that remain consistent across views.
Our study is also limited to a word-level scene-text setting, where lightweight structural screening and string-level consensus already provide useful control signals. Extending this framework to broader OCR settings will require stronger verification than word-level agreement alone. A natural next step is region-level verification, in which candidate transcriptions are tied more directly to localized visual evidence, together with stronger geometric admissibility tests and more precise cross-view validation. We therefore view GRC as a practical first control layer for frozen VLM OCR rather than a complete solution to reliable generative OCR.
6. Conclusion
This paper reframes reliable generative OCR with frozen VLMs as a deployment-control problem rather than an average-accuracy problem. We present a Geometric Risk Controller that places a frozen backbone inside a fixed-protocol selective contract, where multi-view probing provides external evidence, structural screening removes degenerate outputs, and cross-view consensus governs accept-versus-abstain decisions. The result is a small set of predefined operating points that makes the coverage–risk trade-off explicit and auditable.
Across models and datasets, GRC consistently reduces exposed risk while retaining substantial coverage, with the largest gains on long-tail and catastrophic failures. More broadly, the results suggest that for generative perception systems, deployment reliability depends not only on stronger backbones, but also on explicit control over when outputs are exposed to users.