### 15 Apr 2026 v3

# From Plausibility to Verifiability: Risk-Controlled Generative OCR with Vision-Language Models

Abstract.
Modern vision-language models (VLMs) can act as generative optical character recognition (OCR) engines, yet open-ended decoding can expose rare but consequential failures. We identify a core deployment misalignment in generative OCR, where autoregressive decoding favors semantic plausibility while OCR requires outputs that are visually grounded and geometrically verifiable. This mismatch produces severe errors, especially over-generation and unsupported substitutions, creating deployment risk even when benchmark accuracy remains high. We therefore formulate frozen VLM OCR as a selective accept/abstain problem and propose a model-agnostic geometric risk controller. This controller probes multiple structured views of the same input, applies lightweight structural screening, and accepts a transcription only when cross-view consensus and stability satisfy predefined criteria, yielding a small family of operating points. Experiments on frozen VLM backbones and standard OCR benchmarks show consistent reductions in extreme-error risk and catastrophic over-generation at predictable coverage costs. These results suggest that reliable deployment of generative OCR with frozen VLMs benefits from explicit system-level risk control rather than unconstrained generation. Code is available at https://github.com/phare111/GRC.
Generative OCR, Vision-Language Models, Selective Prediction, Risk Control, Abstention
copyright: none
1. Introduction
Figure 1. Layered system view of the proposed geometric risk controller (GRC). The pipeline consists of an input layer, a perception layer, a controller layer under protocol $\pi$, and a selective interface. A frozen vision-language model produces view-wise transcriptions in the perception layer, which are filtered and aggregated in the controller layer to determine whether the system returns a transcription or abstains.
Layered architecture of the geometric risk controller for risk-controlled generative OCR. The figure contains four stages from left to right: an input layer, a perception layer, a controller layer, and a selective interface. An input crop is transformed into \$K\$ geometric views in the perception layer and queried with a frozen vision-language model backbone. The returned transcriptions are then processed in the controller layer through preprocessing, structural screening, cross-view consensus, and operating-point control. The selective interface finally either returns an accepted transcription or abstains with no output.
Optical character recognition (OCR) is increasingly treated as an emergent capability of modern vision-language models (VLMs) (Lv et al., 2024; Yu et al., 2024), which has encouraged the use of frozen VLMs as direct generative OCR engines. However, open-ended autoregressive decoding remains fragile in deployment: models may over-generate beyond the visible text or produce semantically plausible but visually unsupported substitutions. Such failures violate a core requirement of user-facing OCR, namely geometric verifiability, meaning that emitted text should remain locally supportable by image evidence under geometric constraints.
Rather than focusing only on average-case benchmark accuracy (Karatzas et al., 2013), we study deployment risk, that is, the risk exposed on outputs returned to users under a fixed inference protocol. Under this view, frozen-VLM OCR becomes a selective prediction problem with abstention: the system either emits a transcription or abstains, and abstention is part of the deployment contract rather than a convenience feature (Geifman and El-Yaniv, 2017; Wen et al., 2025). This shifts evaluation from a single accuracy number to an auditable risk–coverage trade-off under fixed protocols (Geifman and El-Yaniv, 2017; Angelopoulos et al., 2024).
The severe-error tail is especially sensitive to prompting and decoding. When visual evidence is weak or ambiguous, open-ended generation may fall back on language priors and produce visually unsupported but semantically plausible outputs (Li et al., 2023). As a result, stronger benchmark accuracy does not necessarily reduce user-facing catastrophic failures. Reliability should also not depend on ad hoc prompt or decoding choices, since even small changes can materially alter behavior and undermine reproducibility (Zhao et al., 2021; Rudman et al., 2026; Huang et al., 2024b).
We therefore propose a deployment-centric risk-control framework for generative OCR with frozen VLMs. Its core component, the geometric risk controller (GRC), converts open-ended decoding into a selective system under a fixed accept/abstain contract. Under a fixed multi-view protocol with $K$ geometrically related views, GRC probes the frozen model through controlled input variations, applies lightweight structural validity checks, and accepts a transcription only when cross-view agreement is sufficiently stable. A single strictness knob $m$ exposes a small family of operating points on a risk–coverage curve, enabling explicit operating-point control rather than ad hoc retuning (Geifman and El-Yaniv, 2017; Angelopoulos et al., 2024). We report fixed-protocol results, full component ablations, and compact query-budget analyses so that trade-offs remain comparable across models and datasets. While related in spirit to recent test-time scaling strategies based on augmentation and aggregation (Kaya et al., 2026), our formulation is specialized to OCR failure modes and instantiated as a fixed, deployment-auditable accept/abstain contract.
Our contributions are threefold:
(1) We recast frozen VLM-based OCR as a deployment-control problem, introducing geometric verifiability and deployment-oriented risk primitives that reveal long-tail and catastrophic failures beyond mean accuracy.
(2) We develop a model-agnostic geometric risk controller (GRC), converting open-ended generative OCR into a fixed-protocol selective system with explicit operating points and a controllable strictness knob.
(3) We demonstrate across multiple frozen VLM backbones and scene-text benchmarks that GRC consistently suppresses exposed long-tail risk while providing controllable coverage trade-offs and a compact view of the associated inference-cost trade-off.
2. Related Work
We organize prior work around deployment reliability in frozen VLM OCR, focusing on why stronger OCR capability alone does not guarantee reliable deployment, how prior work introduces selective accept/abstain interfaces to address this gap, and what forms of test-time evidence are needed to make such interfaces auditable under a fixed inference protocol.
2.1. Generative OCR Backbones and Deployment Misalignment
Frozen VLMs as Generative OCR Engines. VLMs are increasingly used as OCR engines (Kim et al., 2022). Systems such as KOSMOS-2.5 (Lv et al., 2024), TextHawk (Yu et al., 2024), and OCEAN-OCR (Chen et al., 2025) show strong end-to-end transcription ability across several settings. However, stronger OCR capability does not by itself guarantee deployment reliability. Under open-ended generation, autoregressive decoding optimizes next-token likelihood rather than explicit visual verifiability, leading to unsupported substitutions and runaway continuations (Huang et al., 2024b). These failures are sensitive to prompting and decoding choices (Rudman et al., 2026; Schmalfuss et al., 2025; Huang et al., 2024b), complicating auditing and control in deployment. Recent analyses further link such behaviors to broader visual-grounding weaknesses in multimodal LLMs (Tong et al., 2024).
Average-Case Metrics versus Deployment Exposure. Benchmarks such as the ICDAR series (Karatzas et al., 2013, 2015; Huang et al., 2019), OCRBench (Liu et al., 2024; Fu et al., 2025), and SVT (Wang et al., 2011) emphasize aggregated scores such as recognition accuracy, CER, and exact match. While these metrics capture average capability, they can understate rare but high-cost tail failures (Amit et al., 2025). For generative OCR, this creates a deployment misalignment between benchmark capability and exposed-output reliability, motivating an explicit accept/abstain formulation (Mao et al., 2025) under selective prediction and risk–coverage principles (Geifman and El-Yaniv, 2017).
$2.$ Selective Deployment and Risk-Controlled Decision Interfaces
Selective deployment requires an explicit accept/abstain interface, not merely a stronger recognizer. Selective prediction formalizes this interface through the coverage–risk trade-off (Geifman and El-Yaniv, 2017; Wen et al., 2025; Chow, 1970), while Conformal Risk Control (CRC) formalizes finite-sample control of expected risk under a specified loss (Wang et al., 2025; Angelopoulos et al., 2024; Angelopoulos and Bates, 2021). In generative systems, however, abstention is often based on internal confidence or self-evaluation (Ren et al., 2023; Mao et al., 2025), and such signals can be prompt- and protocol-sensitive (Schmalfuss et al., 2025) while remaining overconfident on hallucinated outputs (Zollicoffer et al., 2025; Mao et al., 2025; Woo et al., 2025). Our controller adopts the same deployment-level framing but remains empirical rather than guarantee-providing, basing abstention on protocol-induced exogenous evidence—cross-view consistency and structural stability—to define auditable fixed-protocol operating points.
2.3. Verification Signals under Fixed Inference Protocols
Test-time scaling methods such as test-time augmentation, multi-crop inference, and self-consistency are widely used to improve performance and robustness through aggregation, reranking, or voting (Snell et al., 2024; Lee et al., 2025; Wang et al., 2023; Kaya et al., 2026; Chou et al., 2025a, b). Our setting uses repeated inference for a different purpose. Under a fixed protocol, cross-view agreement and stability serve as decision evidence for accept/abstain control rather than as heuristics for improving accuracy (Khan and Fu, 2024). Related OCR and structured extraction systems also use lightweight output constraints, such as character-set, length, and format checks, to suppress obviously invalid generations (Willard and Louf, 2023), but these string-level filters do not directly establish visual support (Li et al., 2025) and can remain misleading when the model is consistently wrong (Huang et al., 2024a). We therefore treat cross-view consistency under controlled geometric perturbations as exogenous evidence of visual support, turning test-time consistency into a fixed-protocol control signal for auditable accept/abstain decisions.
3. Methods
Figure 1 provides an overview of the proposed geometric risk controller (GRC). Given an input crop II, the fixed protocol $\pi$ generates multiple geometric views $\{I_{k}\}_{k=1}^{K}$, which are queried with the frozen generator to produce raw outputs $\{y_{k}\}_{k=1}^{K}$. The controller then canonicalizes these outputs, applies structural screening to obtain the valid set $\Omega(I)$, extracts a consensus candidate $s^{*}(I)$ together with summary statistics such as the vote fraction $q(I)$ and dispersion $\Delta(I)$, and finally applies the operating-point rule to either expose the candidate or abstain.
3.1. Problem Setting and Control Contract
We study generative OCR with a frozen VLM used as a black-box generator. Given an input crop $I\in\mathcal{I}$, a fixed prompt template $p$, and a fixed deterministic decoding configuration $\delta$, the backbone with frozen parameters $\theta$ returns a decoded string
$y=G_{\theta,\delta}(I,p)\in\mathcal{Y},$
where $\mathcal{I}$ is the space of input crops and $\mathcal{Y}$ is the space of decoded strings. Both $\theta$ and $\delta$ are fixed throughout deployment. Thus, $G_{\theta,\delta}$ is treated as a black-box mapping under the fixed prompt-and-decoding setup. Control bypasses model internals, relying only on queried image views and returned strings.
Our goal is not to modify the backbone, but to place it inside a selective accept/abstain contract. Accordingly, the deployed OCR system is a map
$f:\mathcal{I}\to\mathcal{Y}\cup\{\bot\},$
where $\bot$ denotes abstention. The system output is induced by an inference-time evidence protocol $\pi$ together with an acceptance rule applied to the resulting evidence.
Specifically, for an input crop II, protocol $\pi$ produces a candidate transcription $\hat{y}_{\pi}(I)\in\mathcal{Y}$ and an evidence summary $e_{\pi}(I)$. An acceptance rule AA then determines whether the candidate is exposed:
$f_{\pi}(I)=\begin{cases}\hat{y}_{\pi}(I),&\text{if }A\!\left(e_{\pi}(I)\right)=1,\\ \bot,&\text{otherwise.}\end{cases}$
Conventional frozen-VLM OCR is recovered as the degenerate always-accept case, in which the protocol reduces to a single standard query and the returned output is always exposed.
3.2. Evidence Protocol and Multi-View Geometric Probing
We instantiate protocol $\pi$ by querying multiple geometric views of the same input crop II. Let $\mathcal{T}$ denote a family of admissible image transforms, where each $T\in\mathcal{T}$ applies a mild geometric perturbation—such as a small translation, crop jitter, or scale variation—while preserving the nominal text content.
A fixed protocol $\pi$ specifies the transform family $\mathcal{T}$, a rule for selecting concrete transforms, and the total number of queried views $K$. The original crop is always included as the anchor view,
$T_{1}=\mathrm{\textit{Id}},\qquad I_{1}=I,$
where $\mathrm{Id}$ is the identity transform. The remaining views are
$I_{k}=T_{k}(I),\qquad k=2,\dots,K,$
where each $T_{k}\in\mathcal{T}$ is chosen according to the protocol. Thus, II denotes the original input crop, while $I_{k}$ denotes its $k$-th transformed view.
Each view $I_{k}$ is queried with the same prompt template $p$ using the frozen generator $G_{\theta,\delta}$:
$y_{k}=G_{\theta,\delta}(I_{k},p),\qquad k=1,\dots,K.$
Here, $y_{k}$ is the raw decoded string returned for view $I_{k}$. This yields the evidence record
$R_{\pi}(I)=\{(I_{k},y_{k})\}_{k=1}^{K},$
namely, the collection of queried views and their returned strings under protocol $\pi$. Under the fixed protocol, cross-view stability serves as an operational proxy for geometric support, while cross-view disagreement acts as a risk signal for the final accept/abstain decision.
$3.$ Canonicalization and Structural Screening
To make cross-view evidence comparable, we map each raw output $y_{k}$ to a canonical form
$\tilde{y}_{k}=\mathrm{Norm}(y_{k})\in\mathcal{Y},$
where $\mathrm{Norm}(\cdot)$ is a protocol-fixed normalization rule. It removes only presentation variance, such as whitespace normalization and, when required by the task protocol, case canonicalization, without changing the underlying transcription content.
We then apply structural screening through
$v_{k}=\mathbb{I}[\tilde{y}_{k}\in\mathcal{V}]\in\{0,1\},$
where $v_{k}$ is a binary validity indicator and $\mathcal{V}\subseteq\mathcal{Y}$ is a protocol-fixed set of lightweight admissibility constraints. These constraints are label-agnostic and are used only to reject degenerate continuations.
The main constraint in $\mathcal{V}$ is a conservative geometric length bound. For each view $I_{k}$, we compute an upper bound $L_{\mathrm{geom}}(I_{k})$ from foreground geometry after scale normalization and dominant-axis span estimation. Writing $|\tilde{y}_{k}|$ for the character length of $\tilde{y}_{k}$, we mark an output as invalid whenever
$|\tilde{y}_{k}|>L_{\mathrm{geom}}(I_{k}).$
This bound is not intended to estimate the exact ground-truth length. Rather, it serves only to suppress implausibly overlong continuations.
The resulting processed evidence is
$\widetilde{R}_{\pi}(I)=\{(\tilde{y}_{k},v_{k})\}_{k=1}^{K}.$
We then define the valid-view index set
$\Omega(I)=\{k\in\{1,\dots,K\}:v_{k}=1\},$
and its cardinality
$n(I)=|\Omega(I)|.$
All downstream evidence statistics are computed only over the valid set $\Omega(I)$. This induces a minimum-evidence gate with threshold $K_{\min}$.
3.4. Decision Functional and Operating-Point Control
Figure 1 sketches the key quantities formalized below: the consensus candidate $s^{*}(I)$, its support $q(I)$, the dispersion $\Delta(I)$, and the operating-point thresholding rule. For inputs with $n(I)\geq K_{\min}$, let $s^{*}(I)$ denote the unique mode of the valid canonicalized outputs $\{\tilde{y}_{k}:k\in\Omega(I)\}$, when such a mode exists. If no unique mode exists, the controller abstains. Otherwise, the candidate transcription is
$\hat{y}_{\pi}(I)=s^{*}(I).$
To quantify support for the candidate, we define the vote fraction
$q(I)=\frac{1}{n(I)}\sum_{k\in\Omega(I)}\mathbb{I}\!\left[\tilde{y}_{k}=s^{*}(I)\right]\in[0,1],$
that is, the fraction of valid views whose canonicalized outputs exactly match $s^{*}(I)$.
To quantify residual disagreement around the candidate, we define the bounded normalized edit distance
$d(a,b)=\min\!\left(1,\frac{\mathrm{ED}(a,b)}{\max(1,\max(|a|,|b|))}\right)\in[0,1],$
where $d(a,b)$ is the bounded normalized edit distance between strings $a$ and $b$, $\mathrm{ED}(a,b)$ is their Levenshtein edit distance, and $|a|$ denotes the length of string $a$. Using $d(\cdot,\cdot)$, we define the candidate-centered dispersion
$\Delta(I)=\frac{1}{n(I)}\sum_{k\in\Omega(I)}d\!\left(\tilde{y}_{k},s^{*}(I)\right),$
which measures how tightly the valid outputs cluster around the candidate. The resulting evidence summary is
$e_{\pi}(I)=\big(n(I),\,q(I),\,\Delta(I)\big).$
The controller uses two thresholds: a consensus threshold $\tau(m)$ and a dispersion threshold $\kappa$. Here, $m$ is the strictness index of the operating point, and $\tau(m)$ is nondecreasing in $m$. In the main operating-point family, changing $m$ affects only the minimum required vote fraction, while $\kappa$ is held fixed. Larger $m$ therefore induces a stricter acceptance region through a stronger consensus requirement under the common stability gate $\Delta(I)\leq\kappa$.
The instantiated controller under protocol $\pi$ and operating point $m$ is
[…]
where $\bot$ denotes abstention.
Figure 2 provides an intuitive view of this accept/abstain rule under the fixed K=5K{=}5 protocol at the default operating point m=3m{=}3. In the top row, fragmented support leads to abstention because the vote fraction $q(I)$ is too low. In the middle row, a majority candidate still leads to abstention because the valid outputs remain too dispersed, yielding a large $\Delta(I)$. In the bottom row, a borderline majority is accepted because the valid views remain tightly clustered, so both the consensus and stability criteria are satisfied.
Figure 2. Representative accept/abstain evidence patterns under the fixed K=5K{=}5 protocol at the default operating point m=3m{=}3. Fragmented valid outputs cause abstention, a majority can still be rejected under high dispersion, and borderline consensus is accepted when the valid views remain tightly clustered.
A single-column figure with three rows showing evidence patterns under a fixed five-view protocol. In the first row, all five views are valid but disagree, yielding low exact consensus and abstention. In the second row, four views are valid and one is invalid; although a majority mode exists, the remaining valid outputs are highly dispersed, so the sample is rejected. In the third row, all five views are valid and most cluster around a dominant mode, so a borderline vote is still accepted. Each row includes a summary box with the number of valid views, the selected mode, the vote fraction q, the dispersion Delta, and the final decision.
Algorithm 1 summarizes the full inference-time execution of GRC under a fixed protocol/controller pair $(\pi,m)$. Starting from input crop II, it generates $K$ protocol-specified views, queries the frozen generator on each view, canonicalizes the returned strings, applies structural screening, and then evaluates the operating-point rule using $q$ and $\Delta$.
Algorithm 1 Inference-time execution of the Geometric Risk Controller under a fixed protocol
1:Crop II; frozen generator $G_{\theta,\delta}$; prompt $p$; protocol/controller pair $(\pi,m)$
2:Output transcription $s^{*}$ or abstention $\bot$
[…]
4:for $k=1,\dots,K$ do
5:  $\tilde{y}_{k}\leftarrow\mathrm{Norm}(G_{\theta,\delta}(I_{k},p))$,  $v_{k}\leftarrow\mathbb{I}[\tilde{y}_{k}\in\mathcal{V}]$
[…]
7:if $|\Omega|<K_{\min}$ then
8:  return $\bot$
9:Let $s^{*}$ be the unique mode of $\{\tilde{y}_{k}:k\in\Omega\}$
10:if no unique mode exists then
11:  return $\bot$
[…]
13:if $q\geq\tau(m)$ and $\Delta\leq\kappa$ then
14:  return $s^{*}$
15:return $\bot$
Algorithm 1 is an executable summary of Eq. (1). The branch $|\Omega|<K_{\min}$ implements the minimum-evidence gate induced by structural screening, while the final decision branch applies the operating-point rule through $\tau(m)$ and $\kappa$.
3.5. Coverage, Risk, and Operating Points
Selective OCR is evaluated jointly by coverage and conditional risk on exposed outputs. For the controlled system $f_{\pi,m}(I)\in\mathcal{Y}\cup\{\bot\}$, coverage is
$\mathrm{Cov}(\pi,m)=\Pr\!\left[f_{\pi,m}(I)\neq\bot\right].$
If the protocol uses randomized view selection, the probability is taken jointly over the data distribution and the induced protocol randomness.
Let $y^{\mathrm{gt}}$ denote the ground-truth transcription. All controlled-system risk metrics are computed on the covered subset. We report mean character error rate (CER), 99th-percentile CER (P99), and the conditional catastrophic exposure rate
$\mathrm{Meltdown}@\delta(\pi,m)=\Pr\!\left[\mathrm{CER}\!\left(f_{\pi,m}(I),y^{\mathrm{gt}}\right)\geq\delta\ \middle|\ f_{\pi,m}(I)\neq\bot\right],$
with $\delta=2$ in our main setting. Since CER is normalized by ground-truth length, severe over-generation may yield $\mathrm{CER}>1$.
We report results on a small set of predefined operating points $m$ under a common protocol $\pi$. In the main instantiation used in this paper, these operating points correspond to increasing consensus thresholds $\tau(m)$ while the dispersion threshold $\kappa$ is fixed. These settings are fixed before evaluation and shared across models and benchmarks.
For comparison, we report baseline risk on the full test set and controlled-system risk on the covered subset together with coverage. As an additional reference point, we include an approximately coverage-matched external selective baseline in Sec. 4.3.
4. Experiments
4.1. Experimental Setup
Backbones and datasets. We evaluate the controller on three frozen vision-language backbones used as black-box generators: LLaVA-Phi3 (3.8B) (Zhu et al., 2024), Gemma3 (4B) (et al., 2025), and GLM-OCR (Duan et al., 2026). Experiments use the official test sets of IIIT5K (Mishra et al., 2012) and ICDAR 2013 (Karatzas et al., 2013).
Protocol instantiation. Unless otherwise noted, the common protocol uses $K=5$ geometrically related views, consisting of one anchor view and four additional views sampled from a fixed admissible transform distribution over translation, crop jitter, and scale variation. Returned strings are canonicalized by $\mathrm{Norm}(\cdot)$ and filtered by the admissibility screen $\mathcal{V}$, with $K_{\min}=3$.
Operating points. We use predefined strictness levels $m\in\{1,3,5\}$. In the main instantiation, these correspond to increasing consensus thresholds
$\tau(1)=0.1,\qquad\tau(3)=0.5,\qquad\tau(5)=0.9,$
while the dispersion threshold is fixed at $\kappa=0.4$. All settings are fixed before evaluation and shared across models and benchmarks. Unless otherwise stated, the main results use $m=3$.
Metrics and reporting protocol. Following Sec. 3.5, we report coverage together with covered-subset risk, including mean CER, upper-tail CER (P99), and $\mathrm{Meltdown}@2$. Since the baseline never abstains, its risk is reported on the full test set, whereas controlled-system risk is reported on covered outputs. Whenever the controlled system applies output canonicalization, the same canonicalization is also applied to the baseline. Sec. 4.3 further includes an approximately coverage-matched external baseline.
4.2. Main Results
Figure 3. Qualitative OCR cases under the fixed deployment protocol. Each panel shows the crop, ground truth, the always-accept baseline, and the GRC decision at the default operating point $m=3$. Cases include over-generation, unsupported substitution, correct acceptance, cross-view instability, a stable-but-wrong residual failure, and a borderline accepted case. The evidence tag reports valid-view count, consensus fraction $q$, and geometric length estimate $L_{\mathrm{geom}}$; for the instability case, it reports dispersion $\Delta$ instead of $L_{\mathrm{geom}}$.
[…]
Table 1. Main deployment-level results at the fixed operating point ($K=5$, $m=3$).
Dataset Metric Model
LLaVA-Phi3 Gemma3 GLM-OCR
[…]
For each metric, entries are shown as Baseline/System. For coverage, the baseline is always 100.0 because the always-accept system never abstains. For risk metrics, Baseline is computed on the full test set and System on the covered subset. CER is reported in %, MD@2 in ‰.
Table 1 reports the main deployment-level results under the common fixed protocol with $K=5$ at the moderate operating point $m=3$. These results address the deployment setting introduced in Sec. 4.1 by quantifying the error risk exposed on the outputs that the system actually returns under a fixed selective contract.
Across all $3\times 2$ model–dataset settings, the geometric risk controller maintains high coverage while substantially reducing exposed risk relative to the always-accept baseline. Relative to baseline full-set exposure, the controlled system consistently lowers mean CER, suppresses upper-tail error, and reduces catastrophic exposure as measured by $\mathrm{Meltdown}@2$. These results align with the intended deployment role of the controller, turning open-ended generative OCR into a selective system that trades a moderate amount of coverage for a substantially safer set of exposed outputs. Figure 3 shows representative qualitative outcomes under the fixed protocol, including over-generation, unsupported substitution, instability-driven abstention, and stable-but-wrong residual failures.
4.3. Approximately Coverage-Matched External Baseline Comparison
We next test whether the gains of GRC can be explained by standard confidence-based abstention alone. To this end, we compare against an external confidence-threshold selective baseline (Conf.-Thr.) adapted from the vanilla selective-prediction component of Srinivasan et al. (2024). We retain only the threshold-based abstention rule, so that the comparison isolates a standard internal-confidence selective baseline under the OCR setting.
For fairness, this baseline uses the same frozen backbone, prompt template, deterministic decoding, and output canonicalization as the always-accept OCR baseline. It operates on a single view and accepts a transcription only when the model’s internal confidence exceeds a calibrated threshold. Confidence is instantiated by the mean token log-probability of the generated transcription, and the threshold is calibrated on a held-out split to target the coverage of our default operating point $m=3$ for each backbone–dataset setting. Realized test-set coverage is therefore approximately, rather than exactly, matched.
Table 2 reports the comparison. At similar coverage, the confidence-threshold baseline reduces some ordinary errors, but remains much weaker at suppressing catastrophic exposure. GRC achieves lower CER mean and lower $\mathrm{Meltdown}@2$ in all six settings, with the largest gains on the more failure-prone LLaVA-Phi3. This suggests that single-view internal confidence does not capture the instability revealed by multi-view external evidence.
Table 2. Approximately coverage-matched comparison with an external confidence-threshold selective baseline adapted from Srinivasan et al. (2024). For each backbone–dataset setting, the threshold is calibrated on a held-out split to target the coverage of our default operating point $m=3$. Coverage and mean CER are reported in %, Meltdown@2 (MD@2) is reported in ‰.
[…]
IIIT5K LLaVA-Phi3 Conf.-Thr. 90.0 30.9 19.6
Ours 89.5 8.4 0.3
Gemma3 Conf.-Thr. 95.8 7.6 2.2
Ours 94.9 3.7 0.0
GLM-OCR Conf.-Thr. 94.8 4.4 2.2
Ours 95.7 4.0 0.7
ICDAR13 LLaVA-Phi3 Conf.-Thr. 88.2 82.7 31.4
Ours 89.0 9.3 3.7
Gemma3 Conf.-Thr. 95.5 22.5 15.1
Ours 93.0 3.7 1.8
GLM-OCR Conf.-Thr. 95.9 3.2 5.8
Ours 96.2 2.3 0.9
$4.$ Operating-Point Control
Figure 4. Risk–coverage trajectories under the common fixed protocol for all three backbones on IIIT5K and ICDAR13. Each subfigure corresponds to one backbone: (a) GLM-OCR, (b) Gemma3, and (c) LLaVA-Phi3. Each point is a predefined operating point indexed by $m$ while larger $m$ yields lower coverage and lower covered-output risk. Filled markers denote the default operating point $m=3$.
A figure with three horizontally arranged subplots showing risk–coverage trajectories for three OCR backbones: (a) GLM-OCR, (b) Gemma3, and (c) LLaVA-Phi3. In each subplot, the horizontal axis is coverage in percent and the vertical axis is mean character error rate on covered outputs in percent. Two dataset-specific curves are shown, one for ICDAR13 and one for IIIT5K. Each curve consists of predefined operating points indexed by m, with labels marking values such as m=1, m=3, and m=5. Filled markers denote the default operating point m=3. Across all three models, larger m corresponds to lower coverage and lower covered-output risk, with the strongest separation visible for LLaVA-Phi3. A shared legend identifies the two datasets.
Figure 5. Component effects and comparison to an external confidence-threshold baseline on LLaVA-Phi3 under the fixed protocol K=5K{=}5, m=3m{=}3. Bars show Meltdown@2, and labels show coverage. Both structural screening and cross-view consensus reduce catastrophic exposure, while Full GRC achieves the lowest Meltdown@2 on both datasets with competitive coverage.
A single-column figure with two side-by-side bar charts labeled IIIT5K and ICDAR13. Each chart compares five methods—Baseline, Conf-thr, Consensus, Structural, and Full GRC—using Meltdown at 2 on the vertical axis. Numeric labels above the bars report coverage. In both datasets, Full GRC has the lowest catastrophic exposure, while the confidence-threshold baseline reduces exposure relative to the always-accept baseline but remains substantially worse than Full GRC.
A key property of the controller is that the strictness index $m$ does not modify the backbone or protocol. Instead, it selects a predefined operating point under a fixed selective contract. In our main instantiation, increasing $m$ raises the required consensus threshold $\tau(m)$ while the dispersion gate remains fixed, making acceptance more conservative and trading coverage for lower exposed risk. This design makes $m$ a practical deployment knob: the operator does not alter model parameters or inference procedures, but instead selects from a fixed and auditable family of operating points.
Figure 4 visualizes the full risk–coverage trajectories induced by these predefined operating points under the fixed protocol, while Table 3 reports exact numbers for three representative deployment settings, namely the always-accept baseline, a permissive setting $m=1$, and a conservative setting $m=5$. Across all three backbones, increasing strictness moves the system toward lower coverage and lower exposed risk. The figure emphasizes the global trade-off frontier, whereas the table highlights deployment-relevant endpoint statistics, especially catastrophic exposure measured by MD@2. This separation is most pronounced on LLaVA-Phi3 and remains consistent on Gemma3 and GLM-OCR. The default setting $m=3$ used in Sec. 4.2 lies between these endpoints and provides a practical balance between availability and exposed-output reliability.
4.5. Ablation Study
We next ablate the two decision channels of the controller under the same fixed protocol used above. Figure 5 compares the full system with two partial variants and an external confidence-threshold baseline on LLaVA-Phi3. The w/o structural variant removes the protocol-fixed admissibility screen while retaining cross-view consensus control, whereas w/o consensus removes vote-fraction and dispersion control while retaining structural screening.
The two components play distinct and complementary roles. The w/o structural variant reduces catastrophic exposure relative to the always-accept baseline, showing that cross-view consensus alone already provides meaningful risk control. The w/o consensus variant also improves substantially over the baseline, indicating that structural screening is effective at filtering degenerate continuations. However, neither partial variant matches the full controller.
Full GRC achieves the lowest Meltdown@2 on both IIIT5K and ICDAR13 while maintaining competitive coverage. The external confidence-threshold baseline reduces catastrophic exposure compared with the always-accept baseline, but remains clearly worse than Full GRC on both datasets. Taken together, these results show that the controller is not a loose collection of heuristics, but a complementary combination of admissibility screening and cross-view consensus control.
4.6. Effect of Query Budget
Table 3. Representative endpoint operating points under the fixed protocol $K=5$. Coverage (Cov.) is reported in %, and Meltdown@2 (MD@2) is reported in ‰.
IIIT5K ICDAR13
Model Set. Cov. $\uparrow$ MD@2 $\downarrow$ Cov. $\uparrow$ MD@2 $\downarrow$
LLaVA-Phi3 Baseline 100.0 33.7 100.0 35.6
$m=1$ 91.3 0.3 92.1 3.7
$m=5$ 67.7 0.0 68.3 0.9
Gemma3 Baseline 100.0 3.7 100.0 16.4
$m=1$ 95.8 0.0 95.1 1.8
$m=5$ 82.5 0.0 82.0 1.8
GLM-OCR Baseline 100.0 3.7 100.0 1.8
$m=1$ 95.9 1.0 96.5 0.9
$m=5$ 90.5 0.0 91.0 0.9
Table 4 reports a compact query-budget check on LLaVA-Phi3. The single-pass always-accept baseline is included for reference, while $K\in\{3,5,7\}$ varies the evidence budget within the multi-view controller under the same fixed operating protocol. Reducing the number of views to $K=3$ slightly lowers catastrophic exposure, but also lowers coverage, reflecting a more conservative controller with reduced availability. By contrast, increasing the budget to $K=7$ introduces additional query cost with little improvement over $K=5$. Taken together, these results indicate that $K=5$ already captures most of the benefit of multi-view evidence aggregation, making it a practical default for deployment.
Table 4. Compact query-budget check for LLaVA-Phi3 at $m=3$. Relative time is normalized by the single-pass baseline and coverage is reported in %, and Meltdown@2 in ‰.
$K$ Time IIIT5K ICDAR13
Cov. $\uparrow$ MD@2 $\downarrow$ Cov. $\uparrow$ MD@2 $\downarrow$
Baseline 1.0x 100.0 33.7 100.0 35.6
3 2.7x 85.4 0.0 86.3 1.8
5 4.5x 89.5 0.3 88.9 3.7
7 6.3x 89.5 0.3 89.1 3.7
5. Discussion and Limitations
GRC is designed for deployment control rather than backbone improvement. Under a fixed protocol, it reduces unstable user-facing exposure, but it does not eliminate all accepted errors. The main residual failure mode is stable-but-wrong consensus, where multiple admissible views support the same incorrect word and the system still accepts it. This marks a core boundary of the current design. Cross-view agreement provides evidence of stability, but not proof of correctness. As a result, GRC is most effective at filtering unstable and catastrophic exposure, and less effective at detecting errors that remain consistent across views.
Our study is also limited to a word-level scene-text setting, where lightweight structural screening and string-level consensus already provide useful control signals. Extending this framework to broader OCR settings will require stronger verification than word-level agreement alone. A natural next step is region-level verification, in which candidate transcriptions are tied more directly to localized visual evidence, together with stronger geometric admissibility tests and more precise cross-view validation. We therefore view GRC as a practical first control layer for frozen VLM OCR rather than a complete solution to reliable generative OCR.
6. Conclusion
This paper reframes reliable generative OCR with frozen VLMs as a deployment-control problem rather than an average-accuracy problem. We present a Geometric Risk Controller that places a frozen backbone inside a fixed-protocol selective contract, where multi-view probing provides external evidence, structural screening removes degenerate outputs, and cross-view consensus governs accept-versus-abstain decisions. The result is a small set of predefined operating points that makes the coverage–risk trade-off explicit and auditable.
Across models and datasets, GRC consistently reduces exposed risk while retaining substantial coverage, with the largest gains on long-tail and catastrophic failures. More broadly, the results suggest that for generative perception systems, deployment reliability depends not only on stronger backbones, but also on explicit control over when outputs are exposed to users.